Problem 1 (5 points) - Calculate the cross rate between the Costa Rican colon​ (C) and the Canadian dollar​ (C$) from the following spot​ rates: C503.29/$ and C$1.02/$

Problem 2 (10 points) - You are planning a​ 30-day vacation in Bali, Indonesia one year from now. The present charge for a hotel room in Indonesian Rupiah (IDR) is 800,000 IDR/day. The Indonesian Rupiah (IDR) currently trades at IDR16,000/$. You determine that the dollar cost today for a​ 30-day stay would be $1500. The hotel informs you that any increase in its room charges will be limited to any increase in the Indonesian cost of living. Indonesian inflation is expected to be 2.75% per annum, while U.S. inflation is expected to be 1.25%

a. How many dollars might you expect to need one year hence to pay for your​ 30-day vacation?

b. By what percent will the dollar cost have gone​ up?
Answer #1

Problem 1) Cross rate between Costa Rican colon (C) and Canadian dollar (C$) = 503.29/1.02 = C493.422/C$

Problem 2)

a)Calculation of Dollars needed 1 year from today-

Spot rate of year 1= spot rate of year 0*(1+rIDR)/(1+r$)
Where,
r= rate of inflation
Spot rate of year 1= 16000*(1+0.0275)/(1+0.125) 14613.33333
Dollars needed= (30days *(800000IDR/day))/14613.33IDR/$

1642.336141 IDR/$

b) Dollar cost has gone up by (1642.336-1500)/1500= 9.49%